. A random variable X has Poisson distribution with mean equal to 0.4. What is the
probability that the random variable is greater than zero?
1
Expert's answer
2020-06-03T19:11:42-0400

According to the Poisson distribution, the probability of a value to take the value k:

"P(X=k)=\frac{\lambda^k}{k!}e^{-\lambda}"

We know

"E(X)=\lambda=0.4"

Then:

"P(X>0)=1-P(0)=1-e^{-0.4}\approx0.33"